OpenStudy (some.random.cool.kid):

The population of a particular town is given by the function P(t) = 1,200(1.05)2t, where t is the time in years and P(t) is the population after t years. What is the current population, the percentage growth rate, and the population size (rounded to the nearest whole person) in 7 years?

OpenStudy (some.random.cool.kid):

Hint: Percentage Growth Rate = r ⋅ 100 in A = A0(1 + r)t 1,200, 5% annually, and 1,689 1,200, 50% semiannually, and 20,503 1,200, 10.5% per year, and 2,414 1,200, 5% semiannually, and 2,376

zepdrix (zepdrix):

Hey random cooool person. So we have our function,\[\large\rm P(t)=1200(1.05)^{2t}\]

zepdrix (zepdrix):

They gave us a really helpful hint, When the function is in this form:\[\large\rm P(t)=P_o(1+r)^{t}\]Then \(\large\rm 100r\) is our growth rate. Notice, however, that we have a 2 in the exponent of our problem. That's not the form that we want. So we'll have to apply a handy exponent rule to fix that.\[\large\rm (a^b)^c=a^{bc}\]

zepdrix (zepdrix):

So we can use this rule to rewrite our function as,\[\large\rm P(t)=1200\left[(1.05)^2\right]^{t}\]Squaring the 1.05 gives us,\[\large\rm P(t)=1200\left(1.1025\right)^t\]

zepdrix (zepdrix):

Oh they gave us semi-annually for some of the options :( That's really annoying... ugh

OpenStudy (some.random.cool.kid):

back my internet died on me.

zepdrix (zepdrix):

I guess we actually don't need to do all this extra work since we have multiple choice.\[\large\rm P(\color{orangered}{t})=1200(1.05)^{2\color{orangered}{t}}\]We can simply plug in t=7 and simplify,\[\large\rm P(\color{orangered}{7})=1200(1.05)^{2\color{orangered}{\cdot7}}\]You'll need a calculator for this :)

OpenStudy (some.random.cool.kid):

My bad my parents friends just left i had to help escort them, alright so continue :`> whats next? after i calculate it?

zepdrix (zepdrix):

Calculate it :U

zepdrix (zepdrix):

It should match up with one of your options :)

OpenStudy (some.random.cool.kid):

:o oh alright thanks. got it.